An Oil and Gas Company offers a position to Amirul with the starting salary of RM54,000 per year and annual increment of 15%.
(a) Find his salary of fifth years.
(b) Show that his total salary from the first year to the nth years is
RM360,000 (1.15^n - 1)
(c) Find the number of years that he has worked if his total salary is exceed RM3, 654, 670.
1
Expert's answer
2020-12-02T11:47:40-0500

use the compound interest formula :

"a_n = a_0(1+p/100)^n\\
\\ \ \\
a)S_n = 54000(1+0.15)^n\\ 
S_5 = 54000(1+0.15)^5 = 108613.228\\ 
b)from \ 1 \ to \ n \ : \\
54 000 + 54000*1.15 +54000*1.15^2 +... +54000*1.15^n = \\
54000 \sum_0^n(1.15)^n = 54000*1/3(23*1.15^n -20) = 18000(23*1.15^n -20)\\
c) 18000(23*1.15^n -20) = 3654670\\
23*1.15^n -20 = 203.04\\
1.15^n = 9.7\\
n = 16.2571\\
n > 16\\
\text{more than 16 years}"